OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 4 points) A cylindrical rain barrel has a radius of 2 feet and holds a total of 30 cubic feet of water. How tall is the rain barrel? Use 3.14 for pi. Round your answer to the nearest hundredth. 1.58 ft 2.39 ft 3.57 ft 4.78 ft Question 2 (Multiple Choice Worth 4 points) A paper popcorn cone measures 4 inches in diameter and 9 inches high. What is the volume of popcorn that can be contained within this cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 37.68 in3 39.45 in3 40.23 in3 41.90 in3

OpenStudy (anonymous):

@ali2x2 @pooja195 @Peaches15 @phi

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

fan and medal

OpenStudy (michele_laino):

the volume V of a cylinder, whose radius is r and whose height is h, is given by the subsequent formula: \[\Large V = \pi {r^2}h \cong 3.14{r^2}h\]

OpenStudy (michele_laino):

so the measure of the height h, is: \[\Large h = \frac{V}{{3.14{r^2}}} = \frac{{30}}{{3.14 \times {2^2}}} = ...?\]